OpenStudy (anonymous):

Jermoe walked 2/3 mile in the morning and 11/12 in the evening. How many miles did he walk in all? A. 7/12 B. 19/24 C. 1 1/3 D. 1 7/12

OpenStudy (a1234):

2/3 + 11/12 Make them a common denominator. 8/12 + 11/12 19/12, or 1 7/12.

OpenStudy (anonymous):

That's confusing..Can you explain a little more?

OpenStudy (a1234):

Do you know how to multiply fractions?

OpenStudy (anonymous):

A little

OpenStudy (anonymous):

Ok i starting to understan the problem a little more

OpenStudy (a1234):

Okay. The denominator of the second fraction is 12. The denominator of the first fraction is 3. To add fractions, their denominators have to be the same. So, you can multiply the first fraction by 3 to get the denominator of 12 for BOTH the fractions. 3*4 = 12 Also multiply the numerator by the same number you use for the denominator. 2*4 = 8 8/12 + 11/12 11 + 8 = 19 You have 19/12, or 1 7/12.

OpenStudy (anonymous):

So whatever I do to the bottom I do to the top?

OpenStudy (a1234):

Exactly!

OpenStudy (a1234):

But you don't add the denominators (bottom numbers); you keep them the same.

OpenStudy (anonymous):

okok and so it's D then

OpenStudy (a1234):

Yes.

OpenStudy (anonymous):

:) Thank you

OpenStudy (a1234):

You're welcome.